OpenStudy (anonymous):

I need help. I don't understand how to solve this equation. z^2-9z+18=0 Show Steps.

OpenStudy (anonymous):

first factor it (z-6)(z-3)=0 z = 6 or 3

OpenStudy (anonymous):

because either z-6 = 0 or z-3 = 0

OpenStudy (anonymous):

(-b(+-)sqrt(b^2-4ac))/2a. a=w^2 = 1 b=-7w = -7 c=-8 w= (7+-sqrt(49-4*(1*-8)))/2 = (7+-sqrt(81))/2 = (7+-9)/2 = 16/2 and -2/2 => w=8 V w=-1 I wrote this for a similar question. Just insert for a,b and c, and z instead of w, from the equation you provided. V is mathematical notation, meaning "and". So you'll have two answers for z, both equally correct.

OpenStudy (anonymous):

u don't need to do this as the function factorises

OpenStudy (anonymous):

Thanks jimmyrep, I don't understand Msc.student ... Im only in Algebra 1

OpenStudy (anonymous):

Yeah, but factorisation requires some intuition, meaning you'll have to have solved a lot of problems using it, to get the answers right consistently. I think a better place to start is the second degree equation.

OpenStudy (anonymous):

I'm not familiar with how the American school works, so Algebra 1 doesn't mean much to me.

OpenStudy (anonymous):

8th Grade

OpenStudy (anonymous):

Oh, okay. Factorisation it is, then. You'll become familiar with the second degree equation soon enough.